OpenStudy (anonymous):

Can someone PLEASE HELP me create a quadratic equation that cannot be solved by factoring, but can be solved using the quadratic formula? Then identify the values of a, b, and c, and find the solutions using the quadratic formula.

OpenStudy (anonymous):

The most straight-forward way to make an un-factorable equation is to be sure that b is not the sum of any of the factors of c. So, for example, x^2+50x+2 is unfactorable.

OpenStudy (anonymous):

And I guess, not a twice a perfect square, which actually this example seems to be = x^2 + 50x + 625

OpenStudy (anonymous):

So if i used x^2 + 50x + 625... What would the values of a, b, and c be?

OpenStudy (anonymous):

I think telliott means that c should NOT be a perfect square? x^2+50x+625=(x+25)(x+25)

OpenStudy (anonymous):

I'm confused....

OpenStudy (anonymous):

The form of a quadratic is 0=ax^2+bx+c

OpenStudy (anonymous):

I'll just use yours then... if i use x^2+50x+2

OpenStudy (anonymous):

So what would a, b, and c be?

OpenStudy (anonymous):

i don't know how to identify them..

OpenStudy (anonymous):

The number in front of x^2 is always a, the number before x is b, and the number by itself is c.

OpenStudy (anonymous):

Okay.. Well can you help me make an equation to use?

OpenStudy (anonymous):

On the example a=1, b=50, and c=2

OpenStudy (anonymous):

Yeah. What do you want b to be?

OpenStudy (anonymous):

Its okay i'll use the example... So now for the last part. How do i find the solutions using the quadratic formula

OpenStudy (anonymous):

@Ihm ? sorry are you still there

OpenStudy (anonymous):

Ok so the formula is\[x=(-b \pm \sqrt{b ^{2}-4ac})/2a\] First simplify the stuff under the square root. \[ \sqrt{b ^{2}-4ac}\] \[ \sqrt{50 ^{2}-4*1*2}=\sqrt{2492}\] now plug in the rest for each alternative (plus or minus) \[(-50+\sqrt{2492})/2\] \[(-50-\sqrt{2492})/2\] Solve each of these to give you your two solutions.

OpenStudy (anonymous):

Haha, sorry, quadratics are a slow business.

OpenStudy (anonymous):

\[9x ^{2}+15x+7=0\] It can't e solved by simple factoring.. Quadratic formula is needed.. :)

OpenStudy (anonymous):

I still can't get the solutions..

OpenStudy (anonymous):

(robin's equation has imaginary solutions) What's the square root of 2492?

OpenStudy (anonymous):

Wait i'm confused... we're still using the equation we created x^2+50x+2?

OpenStudy (anonymous):

How do i solve that..

OpenStudy (anonymous):

Did you see the post that starts "Ok so the formula is"?

OpenStudy (anonymous):

I might have to go in a few minutes. Picking up from the end of the "ok so the formula is" post: Take the square root of 2492. Now add it to -50 and divide that by 2. That's solution number 1. For solution number 2 do the same thing but subtract root2492 from -50 instead of adding.

OpenStudy (anonymous):

I'm still not getting it..

OpenStudy (anonymous):

@Ihm

OpenStudy (anonymous):

Ok... I"m not sure where you're stuck. See how far you can get with this equation: 0=x^2+7x+8

OpenStudy (anonymous):

I need to use the same equation..

OpenStudy (anonymous):

You need to use the quadratic formula, yes. I want to see where you're stuck. (also I worked out the problem so you just have perfect squares)

OpenStudy (anonymous):

This is the full question: Create your own quadratic equation that cannot be solved by factoring, but can be solved using the quadratic formula. Identify the values of a, b, and c, and find the solutions using the quadratic formula. Show all work to receive credit. What I have answered so far: x^2+50x+2 A= 1 B= 50 C= 2

OpenStudy (anonymous):

I just need the solutions..

OpenStudy (anonymous):

I gave you all the work and all but plugged the numbers into a calculator for you. If you don't want to learn how to do quadratics, ok. If you /do/ want to learn, I'm happy to help. But if you can't even be bothered to plug (-50+root2492)/2 into a calculator, I can't help you.

OpenStudy (anonymous):

No, I want help. I'm just making sure we are still using the same equation

OpenStudy (anonymous):

Went i put it into my calculator it tells me still its in fraction form.. Should my answer be a single number?

OpenStudy (anonymous):

Wait no i just did it again and got -25... @Ihm

OpenStudy (anonymous):

hmm. For this problem you'll get a long decimal. Calculators are funny about order of operations and such. Try plugging in sqrt(2492) hit enter, then add -50 hit enter, then divide by 2.

OpenStudy (anonymous):

I'll be back in a bit. You should get a number between -1 and 0 for this equation.